aubry:

Steve uses an app that shows him how many kilometers he has run to prepare for a marathon. The app said he ran 6.436 kilometers. He wants to post online how many miles he ran. Steve ran ___ miles. (1 mile = 1.609 kilometers)

Pixel:

3.999145 miles or about 4 miles

aubry:

How many ounces of trail mix are in a bag that weighs 0.5675 kilograms? Input only numeric values. (1 pound = 0.454 kg and 1 pound = 16 ounces

Pixel:

1.25112334 pounds

aubry:

If 1 yard = 3 feet and 1 mile = 5,280 feet, how many yards are there in 2 miles?

shasha12:

thanks @aubry did you get your question answered?

aubry:

no

shasha12:

Sorry I don't know how to do this, did you try brainly or an online calculator?

aubry:

brainly gives you wrong answers \

shasha12:

oh ok

Shadow:

@aubry do you know how to do conversions between units

aubry:

not really

Shadow:

1 yard = 3 feet and 1 mile = 5,280 feet, how many yards are there in 2 miles? So we are given 2 miles, and we need to convert them into yards \[\frac{ 2miles }{ 1} \times \frac{ 5280feet}{ 1mile } \times \frac{ 1yard }{ 3feet }\]

Shadow:

That is how you would solve for it, as when you multiply across \[\frac{ mile }{ mile } = 1\] \[\frac{ feet }{ feet } = 1\] Effectively canceling out. Then you are left with yards.

aubry:

so the answer will be what

Shadow:

Multiply across

Shadow:

\[\frac{ 2 \times 5280 }{ 3 }\]

aubry:

=10560

Shadow:

Yes then divide by 3